myininaya (myininaya):

I think I'm throwing in the towel on this one. \[\int \frac{x^5-x}{x^8+1} dx\] I know @amistre64 can do it though. :p This is @ganeshie8 's evil problem.

myininaya (myininaya):

I have wrote this evil thing in several different ways...

myininaya (myininaya):

\[\frac{x(x^4-1)}{x^8+1}=\frac{x(x^4-1)}{(x^4-\sqrt{2}x^2+1)(x^4+\sqrt{2}x^2+1)}\] I was thinking this form might help because I peaked at the answer on wolfram. http://www.wolframalpha.com/input/?i=int%28%28x%5E5-x%29%2F%28x%5E8%2B1%29%2Cx%29

OpenStudy (amistre64):

i did that when i was 3 ... ;)

myininaya (myininaya):

partial fractions looks ugly

myininaya (myininaya):

I was hoping for a trick. I knew you did it at 3 so I knew you would know all the tricks to this one.

OpenStudy (amistre64):

-x+x^5-x^8+- .... --------- 1+x^8 | -x+x^5 x-x^8 -------- x^5 - x^8 -x^5 -x^13 ---------- -x^8-x^13

OpenStudy (amistre64):

thats my approach, power series representation

OpenStudy (amistre64):

im sure hartnn could whip up a laplace transform to put us all to shame :)

ganeshie8 (ganeshie8):

@hartnn

OpenStudy (amistre64):

my power series approach migh tlook better with a split fraction ... same same tho

ganeshie8 (ganeshie8):

rational function.. so i think many ways to approach... myininaya's first method looks more straightforward to me.. :)

myininaya (myininaya):

I don't like that way though. It is not cute at all.

OpenStudy (dumbcow):

haha cute :)

OpenStudy (amistre64):

i hear fourier series are a god send as well .. cant say ive had much practice with them tho

myininaya (myininaya):

I was looking to fall in love with this integral. I tried to get to know it and it does something disgusting.

OpenStudy (amistre64):

math is evil :) a fickle thing she is

myininaya (myininaya):

ganeshie8 said he loves this problem or likes So I was just looking for the same relationship.

ganeshie8 (ganeshie8):

dividing x^4 top and bottom may bring back your good feelings i hope : \[\int \frac{x^5-x}{x^8+1} dx = \int \dfrac{x-\frac{1}{x^3}}{x^4+\frac{1}{x^4}}dx\]

ganeshie8 (ganeshie8):

thats only half of the trick though.. the other half is in packing the denominator in an useful way..

ganeshie8 (ganeshie8):

noticing that derivative of 1/x^2 is -1/x^3

myininaya (myininaya):

Yeah that is one of the ways I wrote it. In fact I divided by everything (at different times of course). I will tried to play with this one in particular more though.

ganeshie8 (ganeshie8):

\[\large x^4 + \dfrac{1}{x^4} = \left( x^2+\dfrac{1}{x^2}\right)^2 - 2\]

Miracrown (miracrown):

well, moving the troublesome part into the numerator, usually helps. i'd be tempted to try a u-sub with u=x^8+1

myininaya (myininaya):

no more i see gosh neat i do love it again

myininaya (myininaya):

well yes more if people have other ways

myininaya (myininaya):

But I see @ganeshie8 's way now

myininaya (myininaya):

I really think I would have never thought of that

ganeshie8 (ganeshie8):

somebody in openstudy only showed me this solution... i think its eliassaab

ganeshie8 (ganeshie8):

but i feel there should be other more obvious and simple ways to work this...

myininaya (myininaya):

I don't know about that.

myininaya (myininaya):

However I could be wrong.

myininaya (myininaya):

@SithsAndGiggles Always has tricks up his sleeves too.

OpenStudy (perl):

wolfram has steps to solve integrals like this

OpenStudy (perl):

but i think you have to subscribe

OpenStudy (perl):

wolfram *shows* steps

myininaya (myininaya):

And yeah I don't want to subscribe I need my money

OpenStudy (anonymous):

A substitution chain could make the original integral look a bit cleaner. \[\begin{align*}\int\frac{x^5-x}{x^8+1}~dx&=\int\frac{x^5-x}{\left(x^4\right)^2+1}~dx\\\\ &=\int\frac{t^{5/4}-t^{1/4}}{t^2+1}\left(\frac{1}{4}t^{-3}{4}~dt\right)&\text{setting }t=x^4\\\\ &=\frac{1}{4}\int\frac{t^{2/4}-t^{-2/4}}{t^2+1}~dt\\\\ &=\frac{1}{4}\int\frac{t-1}{t^{1/2}\left(t^2+1\right)}~dt\\\\ &=\frac{1}{4}\int\frac{y^2-1}{y\left(y^4+1\right)}(2y~dy)&\text{setting }y=\sqrt t\\\\ &=\frac{1}{2}\int\frac{y^2-1}{y^4+1}~dy \end{align*}\]

OpenStudy (anonymous):

Second line should have \(\large\color{red}{t^{-3/4}}\)

OpenStudy (dan815):

\[\int\]

OpenStudy (dan815):

=]

myininaya (myininaya):

Thank you all for your comments. I'm closing this but feel free to add anything you want. (I mean on this integral or other similar integrals)

OpenStudy (ikram002p):

its easy :P